\begin{thebibliography}{ABC}
%
\bibitem{BR} {\sc Berstel J., Reutenauer C.}, {\it Rational series and their 
languages EATCS Monographs on Theoretical Computer Science}, Springer (1988).\\
New electronic version\\ 
{\tt http://www-igm.univ-mlv.fr/~berstel/LivreSeries/LivreSeries.html}
%
\bibitem{BBM} {\sc Bender C. M., Brody D. C., and Meister B. K.}, 
Quantum field theory of partitions, J. Math. Phys. Vol 40 (1999)
%
\bibitem{GOF7} {\sc Blasiak P., Horzela A., Penson K. A., Duchamp G. H. E. , Solomon A. I.}, 
{\it Boson normal ordering via substitutions and Sheffer-Type Polynomials}, 
Phys. Lett. A {\bf 338} (2005) 108
%
\bibitem{GOF8} {\sc Blasiak P., Penson K. A., Solomon A. I.,  
A. Horzela, G. H. E. Duchamp}, 
{\it Some useful formula for bosonic operators}, Jour. Math. Phys.
 {\bf 46} 052110 (2005).
%
%\bibitem{B_Alg_III} {\sc Bourbaki N.}, {\it Algebra, chapter III}, Springer
%
%\bibitem{B_Alg_VI} {\sc Bourbaki N.}, {\it Algebra, chapter VI}, Springer
%
%\bibitem{B_ST} {\sc Bourbaki N.}, {\it Theory of sets}, Springer
%
%\bibitem{Ca} {\sc Cartier P.}, {\it S\'eminaire ``Sophus Lie'', 2\`eme ann\'ee}, Facult\'e des Sciences de Paris (1955-56)
%
%\bibitem{Ca2} \textsc{Cartier P.}, {\it Fonctions polylogarithmes, nombres polyzeta et groupes pro-unipotents }, S\'eminaire Bourbaki, Asterisque n. %282 (2002) 
%
%\bibitem{CP} {\sc Chari V., Pressley A.}, A guide to quantum groups. 
%Cambridge (1994).
%
%\bibitem{Co1} {\sc Comtet L.}, Analyse Combinatoire, PUF, Tome 1-2 (1970)
%
\bibitem{GOF4} {\sc Duchamp G. H. E., Blasiak P., Horzela A., Penson K. A., Solomon A. I.}, Feynman graphs and related Hopf algebras, Journal of Physics: Conference Series, SSPCM'05, Myczkowce, Poland.
arXiv : cs.SC/0510041
%
%\bibitem{DFLL} {\sc Duchamp G., Flouret M., Laugerotte E., Luque J-G.}, 
%{\it Direct and dual laws for automata with multiplicities}\\
%{\tt arXiv} : math.CO\/0607412
%
%\bibitem{DL} {\sc Duchamp G., Luque J-G.}, 
%{\it Congruences Compatible with the Shuffle Product}\\
%{\tt arXiv} : math.CO\/0607419 
%
%\bibitem{DHT} {\sc Duchamp G., Hivert F., Thibon J. Y.}, {\it Non commutative 
%functions VI: Free quasi-symmetric functions and related algebras}, 
%International Journal of Algebra and Computation Vol 12, No 5 (2002).
%
%\bibitem{DKPTT} {\sc Duchamp G. H. E., Koshevoy G., Penson K. A., Tollu C., Toumazet F.}, {\it Geometric combinatorial twisting and shifting}, %S\'eminaire Lotharingien (in preparation).
%
\bibitem{FPSAC07} {\sc Duchamp G. H. E., Luque J. -G., Novelli J. -C., Tollu  C., Toumazet F.}, {\it Hopf algebras of diagrams}, submitted to FPSAC07.
%
\bibitem{OPG} {\sc Duchamp G., Solomon A. I., Penson K. A., Horzela A., Blasiak P.}, {\it One-parameter groups and combinatorial physics}, 
Proceedings of the Symposium Third International Workshop on
Contemporary Problems in Mathematical Physics (COPROMAPH3)
(Porto-Novo, Benin, Nov. 2003),  J. Govaerts, M. N. Hounkonnou and A. Z. Msezane (eds.), p.436 (World Scientific Publishing 2004)\\
arXiv: {\tt quant-ph/04011262}
%
\bibitem{D21} {\sc Duchamp G., Reutenauer C.}, {\it Un crit\`ere de rationalit\'e 
provenant de la g\'eom\'etrie noncommutative} Invent. Math. {\bf 128} 613-622.
(1997).
%
%\bibitem{F1} {\sc Flajolet P.}, http://algo.inria.fr/flajolet/
%
%\bibitem{Fo1} {\sc Foissy L.}, {\it Isomorphisme entre l'alg\`ebre des fonctions quasi-sym\'etriques libres et une alg\`ebre de Hopf des arbres enracin\'es d\'ecor\'es plans}, personal communication.
%
%\bibitem{Fo2} {\sc Foissy L.}, {\it Les alg\`ebres de Hopf des arbres enracin\'es decor\'es}, PhD Memoir, Reims University (2002). 
%
%\bibitem{Ha} {\sc Hazewinkel M.}, {Hopf algebras of endomorphisms of Hopf algebras}, (Oct 2004) 
%{\tt ArXiv : math.QA/0410364}
%
%\bibitem{Ho} {\sc Hoffman M. E.}, {\it Quasi-shuffle products}, J. Algebraic Combin. (2000), 49-68
%
\bibitem{GOF6} {\sc Horzela A., Blasiak  P., Duchamp  G. H. E., Penson K. A. and Solomon A.I.}, {\it A product formula and combinatorial field theory}, 
Proceedings of the XI International Conference on Symmetry Methods
in Physics (SYMPHYS-11) (Prague, Czech Republic, June 2004), C. Burdik, O. Navratil, and S. Posta (eds.) (JINR Publishers, Dubna)\\ 
arXiv:{\tt quant-ph/0409152}
%
\bibitem{Ha} {\sc Hadamard J.}, {\it Théorème sur les séries entières}, Acta Math., Uppsala, t. 22, 1899, p. 55-63.
%
%\bibitem{JR} {\sc Joni S. A., Rota G.-C.}, {\it Colgebras and Bialgebras in Combinatorics}, Studies in Applied Mathematics {\bf 61}, 93-139 (1979).
%
\bibitem{Ju} {\sc Jungen R.} {Sur les séries de Taylor n'ayant que des singularités algébrico-logarithmiques sur leur cercle de convergence}, Comment. math. Helvet., t. 3 1931, p. 266-306 (Thèse Sc. math.). 
%
\bibitem{D20} {\sc Katriel J., Duchamp G.}, 
{\it Ordering relations for q-boson operators, continued fractions techniques, 
and the q-CBH enigma}. Journal of Physics A {\bf 28} 7209-7225 (1995).
%
%\bibitem{K1} {\sc Knuth D.}, {\it The art of computer programming} Tome I.
%Addison-Wesley (1981)
%
%\bibitem{MM} {\sc Milnor  J., Moore J.}, {\it On the structure of Hopf Algebras}, Ann. of Math. 81 (1965), 211-264. 
%
%\bibitem{Oc} {\sc Ochsenschl\"ager P.}, Binomialkoeffitzenten und Shuffle-Zahlen, Technischer
%Bericht, Fachbereich Informatik, T. H. Darmstadt,1981.
%
%\bibitem{Re} {\sc Reutenauer C.}, {\it Free Lie algebras}, Oxford University Press (1993).
%
%\bibitem{Ro} {\sc Rosso M.}, {\it Quantum groups and quantum shuffles}, Inventiones Math. 133 (1998).
%
\bibitem{Sc1} {\sc Sch\"utzenberger M. P.}, On a theorem of R. Jungen, Proc. American Math. Society (1962) p. 189-197.
%
\bibitem{GOF2} {\sc Solomon A.I., Blasiak P., Duchamp  G. H. E., Horzela A. and Penson K. A.}, {\it Combinatorial Physics, 
Normal Order and Model Feynman Graphs}, Proceedings of the Symposium 'Symmetries in Science XIII', Bregenz, 
Austria, 2003, B. Gruber, G. Marmo and N. Yoshinaga (eds.), p.527 (Kluwer Academic
Publishers 2004)\\
arXiv : {\tt quant-ph/0310174}
%
\bibitem{GOF3} {\sc Solomon A.I., Duchamp G. H. E., Blasiak  P., Horzela  A. and Penson K. A.}, 
{\it Normal Order: Combinatorial Graphs Quantum Theory and Symmetries}, Proceedings of 
the 3rd International Symposium
P.C. Argyres, T.J. Hodges, F. Mansouri, J.J. Scanio, P. Suranyi, and L.C.R.
Wijewardhana (eds.), p.398 (World Scientific Publishing 2004)\\
arXiv:{\tt quant-ph/0402082}
%
\bibitem{GOF5} {\sc Solomon A.I., Duchamp G. H. E., Blasiak P., Horzela A. and Penson K. A.}, {\it Partition functions and graphs: A combinatorial approach}, 
Proceedings of the XI International Conference on Symmetry Methods
in Physics (SYMPHYS-11) (Prague, Czech Republic, June 2004), C. Burdik, O. Navratil, and S. Posta (eds.) (JINR Publishers, Dubna, 2004)\\
arXiv:{\tt quant-ph/0409082}
%
%\bibitem{St2} {\sc Stanley R. P.}, {\it Enumerative Combinatorics}, vol. {\bf 2}, Cambridge Univ. Press (2004).
%
\bibitem{Varvak} {\sc Varvak  A.}, {\it Rook numbers and the normal ordering problem}, J. Comb. Theory, Ser. A 112(2): 292-307 (2005)
%
\end{thebibliography}

\documentclass[12pt]{article}

\usepackage[dvips]{epsfig}
\usepackage{amsfonts}
\usepackage{amscd}
\usepackage{amsmath}
\usepackage{pstricks}
 \font\tensym=msbm10
 \font\sevensym=msbm7
 \font\fivesym=msbm5

 \font\tengoth=eufb10
 \font\sevengoth=eufb7
 \font\fivegoth=eufb5

\newfam\symfam
\textfont\symfam=\tensym
\scriptfont\symfam=\sevensym
\scriptscriptfont\symfam=\fivesym

\newfam\gothfam
\textfont\gothfam=\tengoth
\scriptfont\gothfam=\sevengoth
\scriptscriptfont\gothfam=\fivegoth

\def\hs{\hbox to 3mm{}}
\def\hhs{\hbox to 5cm{}}

\def\JPicScale{0.8}\ifx\JPicScale\undefined\def\JPicScale{1}\fi

\def\C{\mathbb{C}}

\def\N{\mathbb{N}}
\def\Z{\mathbb{Z}}

\def\H{\mathcal{H}}

\def\HWS{{\cal H}_{WS}}
\def\1H{\mathbf{1}_{\HWS}}

\def\al{\alpha}
\def\be{\beta}

\def\2m#1#2{(#2 #1)}
\def\3m#1#2#3{(#3 #2 #1)}

\def\ncp#1#2{#1\langle #2\rangle}

\def\scal#1#2{\langle #1 | #2 \rangle}

\def\ra{\rightarrow}

\def\adots{\mathinner{\mkern2mu\raise1pt\hbox{.}
\mkern3mu\raise4pt\hbox{.}\mkern1mu\raise7pt\hbox{.}}}
\def\binomial#1#2{\begin{pmatrix}#1 \\ #2\end{pmatrix}}

\def\up#1{\raise 1ex\hbox{\footnotesize#1}}

\def\mref#1{{\footnotesize ({\ref{#1}})}}

\newtheorem{expl}{Example}[section]

\newtheorem{theorem}[expl]{Theorem}
\newtheorem{corollary}[expl]{Corollary}

\newtheorem{proposition}[expl]{Proposition}

\begingroup
\def\2#1{\ifnum#1<10 0\fi\the#1}
\xdef\isodayandtime{
{\2\day-\2\month-\the\year\space\2{\count0}:%
\2{\count2}}}
\endgroup

\title{\Large\bf Rational Hadamard products via\\ Quantum Diagonal Operators}

\author{{\sc G. H. E. Duchamp, S. Goodenough}   
\rm\thanks{LIPN - UMR 7030
CNRS - Université Paris 13
F-93430 Villetaneuse, France}\\
{\rm and }{\sc K. A. Penson}
\rm\thanks{Laboratoire de Physique Th\'eorique de la Mati\`ere Condens\'ee, 
Universit\'e Pierre et Marie Curie, CNRS UMR 7600
Tour 24 - 2i\`eme \'et., 4 pl. Jussieu, F 75252 Paris Cedex 05, France}           
}
\date{}
\begin{document}

\maketitle
%\isodayandtime

\begin{abstract} We use the remark that, through Bargmann-Fock representation, diagonal operators of the Heisenberg-Weyl algebra are scalars for the Hadamard product to give some properties (like the stability of periodic fonctions) of the Hadamard product by a rational fraction. In particular, we provide through this way explicit formulas for the multiplication table of the Hadamard product in the algebra of rational functions in $\C[[z]]$.\footnote[0]{Partially supported by the ANR CombPhys.}
\end{abstract}

\section{Heisenberg-Weyl algebra and Bargmann-Fock\\ representation}

Let $\H=l^2(\N,\C)$, the standard separable Hilbert space of complex-valued sequences $(\al_n)_{n\in \N}$ such that $\sum_{n\in\N}|\al_n|^2<\infty$ and $e_n=(\delta_{m,n})_{m\in \N}$ be its the canonical Hilbert basis with 
\begin{equation}
\H_0=\bigoplus_{n\in \N} \C e_n	
\end{equation}
the standard dense subspace of $\H$.\\
Using the classical {\it bra} and {\it ket} denotations, one defines the following operators 
\begin{equation}
	a=\sum_{n\geq 1}\sqrt{n}\ |n-1\rangle\langle n|\ ;\ a^\dag=\sum_{n\geq 0}\sqrt{n+1}\ |n+1\rangle\langle n|
\end{equation}
where ``$|n\rangle$'' (resp. ``$\langle n|$'') stand for ``$|e_n\rangle$'' (resp. ``$\langle e_n|$'').
By definition, $a,a^\dag$ are hermitian conjugates and densely defined (because $\H_0$ is a subspace of $dom(a)$ and $dom(a^\dag)$) and one verifies easily that $[a,a^\dag]=Id_{\H_0}$. Due to their proeminent importance in quantum mechanics, there is a huge litterature on these operators and it can be proved that 
\begin{itemize}
	\item the algebra (in $End(\H_0)$) generated by $a,a^\dag$, can be presented as 
$$
HW_\C=\Big< a,a^\dag\ ;\ aa^+-a^+a=1\Big>_{\C-AAU}
$$
where $\C-AAU$ is the category of $\C$-associative algebras with unit. 	
	\item there is no non-zero representation of $HW_\C$ in a Banach algebra
	\item the representation $\be$ on the space of complex formal power series $\C[[z]]$ by operators $\be(a),\be(a^\dag)$ such that $\be(a)(S)=\frac{d}{dz}S\ ;\ \be(a^\dag)(S)=zS$, known as Bargmann-Fock representation,  is faithful
	\item due to the (only) relator $aa^+-a^+a=1$, the family 
	$$\Big((a^\dag)^k a^l\Big)_{k,\, l\geq\, 0}$$
	 is a basis of $HW_\C$ (the expressions w. r. t. this basis are called ``normally ordered'')
	\item denoting $HW^{(e)}_\C$ the subspace generated by the ``monomials'' $\Big((a^\dag)^k a^l\Big)_{k-l=e}$, one has
\begin{equation}\label{gr_alg}
	HW^{(e_1)}_\C.HW^{(e_2)}_\C\subset HW^{(e_1+e_2)}_\C 
\end{equation}
so that the algebra $HW_\C$ is $\Z$-graded (the parameter $e$ is called ``degree'' by algebraists and ``excess/defect'' -according to its sign - by physicists). 
\end{itemize}

Despite of its simple definition $HW_\C$ supports very rich combinatorial studies \cite{BBM,GOF7,GOF8,GOF4,FPSAC07,OPG,GOF6,D20,GOF2,GOF3,GOF5,Varvak}.

\section{Problem of the rational Hadamard table}

The Hadamard product on generating functions was introduced by the French mathematician Jacques Hadamard \cite{Ha} as a shifted convolution product on the one-dimensional torus ($S^1$, the commutative group of angles). This product (denoted $\odot$ in the sequel) amounts to performing the pointwise product on the coefficients of the expanded sequences 
\begin{equation}
	\Big(\sum_{n=0}^\infty a_n z^n\Big)\odot \Big(\sum_{n=0}^\infty b_n z^n\Big):=\Big(\sum_{n=0}^\infty a_nb_n z^n\Big).
\end{equation}
In \cite{Ju}, it is proved that the Hadamard product of an algebraic and a rational series is algebraic (this theorem was extended later to the noncommutative case by Sch\"utzenberger \cite{Sc1} and the latter could be used as a crucial result in \cite{D21}). Hence the algebraic series are a module over the algebra of rational series and it remains the combinatorial problem of giving explicit formulas for the multiplication of two rational series i. e. elements of $K[[z]]$ of the form $\frac{P(z)}{Q(z)}$ with $Q(0)\not=0$. This problem is called here the ``Rational Hadamard product problem''. We can make this problem very precise in the case when the coefficients are taken in an algebraically closed field (for the sake of readability, it will be here taken equal to $\C$, the field of complex numbers) as the algebra of rational series admits 
\begin{equation}\label{rat_lin_basis}
	(z^n)_{n\in \N} ; \Big(\frac{1}{(1-\al z)^m}\Big)_{\al\in \C^*\atop m\in \N^*}
\end{equation}
as linear basis.
  
The aim of this paper is to show how to compute explicitely the multiplication table of this algebra with respect to the former basis using the classical creation/annihilation operators. To this end, we will use the algebra generated by the two operators on $\C[[z]]$ 
\begin{equation}
\be(a^+):	S\ra zS\ ;\ \be(a):	S\ra \frac{d}{dz} S\ .
\end{equation}
These formulas define (on the vector space of complex series) a faithful representation of the Heisenberg-Weyl algebra $HW_\C$.\\
It follows from Eq \mref{gr_alg} that $HW^{(0)}_\C$ is a subalgebra of $HW_\C$, called the algebra of diagonal operators. In order to compute the Hadamard products of the elements of the basis \mref{rat_lin_basis}, one remarks that the diagonal operators are scalars for the Hadamard product. Indeed, as $HW^{(0)}_\C$ is the linear span of the monomials $(a^+)^na^n$, it is sufficient to prove that, for all $S,T\in \C[[z]]$, 
\begin{equation}
\be((a^+)^na^n)[S]\odot T=S\odot \be((a^+)^na^n)[T]	
\end{equation}

which is easily seen by direct computation 
\begin{eqnarray}
\be((a^+)^na^n)[S]\odot T&=& \sum_{m=0}^\infty \scal{\be((a^+)^na^n)[S]}{z^m}\scal{T}{z^m}z^m= 
\sum_{m=0}^\infty \frac{n!}{m!}\scal{S}{z^m}\scal{T}{z^m}z^m\cr
&=&\sum_{m=0}^\infty \scal{S}{z^m}\scal{\be((a^+)^na^n)[T]}{z^m}z^m=S\odot \be((a^+)^na^n)[T]	
\end{eqnarray}

\section{Results}

In this section, we prove that the Hadamard multiplication operator 

\begin{equation}
f\mapsto \frac{1}{(1-z)^{k+1}}\odot f	
\end{equation}
is a diagonal operator and hence, the corresponding multiplication operator by $\frac{1}{(1-\al z)^{k+1}}$ is the composition of a diagonal operator and a dilation (i. e. a substitution $z\mapsto \al z$) which is itself a Hadamard multiplication operator as 
\begin{equation}
	f(\al z)=\frac{1}{1-\al z}\odot f(z)\ .
\end{equation}

\begin{proposition}
For $k\in\N,\ \al\in \C^*$, let $D_{(\al,k)}$ be the operator ``Hadamard multiplication by 
$\frac{1}{(1-\al z)^{k+1}}$''  
\begin{equation}
f\mapsto \frac{1}{(1-\al z)^{k+1}}\odot f	
\end{equation}
Then, one has \\
1) i) $D_{(1,k)}$ is a diagonal operator through the Bargmann-Fock repesentation $\beta$. More precisely 
\begin{equation}
D_{(1,k)}=\be\Big(\sum_{j=0}^k \frac{\binomial{k}{j}}{j!} (a^+)^ja^j\Big)	
\end{equation}
ii) $D_{(\al,0)}$ is the substitution $z\ra \al z$ (automorphism for $\al\not=0$).\\
iii) Due to the associativity of the Hadamard product, one has 
\begin{equation}
D_{(\al,k)}=D_{(\al,0)}\circ D_{(1,k)}=D_{(1,k)}\circ	D_{(\al,0)}	
\end{equation}
\end{proposition}

These formulas on diagonal operators yield following theorem.

\begin{theorem} 
The multiplication table of the algebra of rational power series is the following
\begin{equation}\label{mult_by_monomials}
z^n\odot z^m=\delta_{m,n}z^n\ ;\ z^n\odot \frac{1}{(1-\al z)^{m+1}}=\binomial{-(m+1)}{n}\al^nz^n
\end{equation}
and 
\begin{eqnarray}\label{mult_by_inverses}
\frac{1}{(1-\al z)^{k+1}}\odot \frac{1}{(1-\al z)^{l+1}}=
D_{(\al\be,0)}\Big[\sum_{j=0}^{k} \frac{\binomial{k}{j}}{j!} (l+1)...(l+j)\frac{z^j}{(1-z)^{l+j+1}}\Big]=\cr
D_{(\al\be,0)}\Big[\sum_{j=0}^{k} \frac{\binomial{k}{j}}{j!} (l+1)^{\bar j} \sum_{s=0}^j \binomial{j}{s} \frac{(-1)^{j-s}}{(1-z)^{l+s+1}}\Big]	
\end{eqnarray}
\end{theorem}

\begin{corollary}
The algebra of rational power series $\ncp{\C}{z}^\circ$ is closed under the action of $HW_\C$
\end{corollary}

\section{Concluding remarks}

In a theorem of G. Cauchon (reproved by other means in \cite{D21}, see also \cite{BR}), one has to consider the algebra of rational series in the sense of Neumann-Malcev algebras. In one variable, it is the field of fractions
\begin{equation}
	\frac{P(z)}{Q(z)}\ ;\ P,Q\in \C[z],\ Q\not=0
\end{equation}
within $\C((z,z^1))$, the algebra of Laurent series. This algebra being a subfield of functions $\Z\ra \C$, the Hadamard product (which is in fact the pointwise product) extends at once.\\
It is immediate to see that this field of frations admits the linear basis 
\begin{equation}\label{full_rat_lin_basis}
	(z^n)_{n\in \Z}\ ; \Big(\frac{1}{(1-\al z)^m}\Big)_{\al\in \C^*\atop m\in \N^*}
\end{equation}
and the multiplication table (formulas \mref{mult_by_monomials} and \mref{mult_by_inverses}) have to be extended by zero as regards the products 
$z^n\odot \frac{1}{(1-\al z)^{m+1}}$ with $n<0$. As a consequence, this algebra again is closed under Hadamard products.

\end{document}